OpenStudy (anonymous):

Help with graphing r=4Sin(5 [theta])

zepdrix (zepdrix):

\(\Large\rm r=4\sin(5\theta)\) Hey Adam! So what's the plan here? We need to construct a polar graph of this?

OpenStudy (anonymous):

yes please :D

OpenStudy (dangerousjesse):

Wolfram can be fairly useful at such times :P

zepdrix (zepdrix):

Umm ok we're going to get some kind of Rose Curve. \(\Large\rm r=a\sin(n\theta)\) \(\Large\rm a\) tells us how far out each petal will extend. \(\Large\rm n\) is telling us how many petals we'll have. (5 in this case). I like to construct a table I guess.. usually helps with plotting it.

OpenStudy (anonymous):

yeah i figured it would form the shape of a rose petal

OpenStudy (anonymous):

anyone here?

OpenStudy (anonymous):

@DangerousJesse thanks!!! i tried using that site, it didnt work for me :(

zepdrix (zepdrix):

Ah sorry D: Had to get food out of the oven.

zepdrix (zepdrix):



zepdrix (zepdrix):

A rose with an `even` number of petals will be traced out in \(\Large\rm 2pi\). Since we have a rose with an `odd` number of petals, it will be traced out in \(\Large\rm \pi\). So let's ummm... choose some angles...

zepdrix (zepdrix):

Desmos is also a nice graphing tool that you can use :) https://www.desmos.com/calculator/kx8ewlwhom

OpenStudy (anonymous):

wow, i wish i would've found out about these sites a long time ago

zepdrix (zepdrix):

So it looks like it's finishing each petal in pi/5. Ugh these are tough to come up with good angles >.<